# Sensor-Based Distributionally Robust Control for Safe Robot Navigation in Dynamic Environments

### v1

@@ 4 Problem Formulation
We consider a mobile robot that relies on noisy range measurements to traverse an unknown dynamic environment towards a desired goal. The environment contains both static and dynamic obstacles. We define the obstacle space as a closed set ${\cal O}(t)\subset\mathbb{R}^{p}$, and the free space as an open set ${\cal F}(t)=\mathbb{R}^{p}\setminus{\cal O}(t)$.
The robot’s motion is governed by control-affine dynamics as in (1). Let $\phi:{\cal X}\rightarrow\mathbb{R}^{p}$ project the robot state $\mathbf{x}$ to its position $\phi(\mathbf{x})\in\mathbb{R}^{p}$, and denote the robot’s body as ${\cal B}(\mathbf{x})\subset\mathbb{R}^{p}$.
@@
The robot is equipped with a range sensor (e.g., LiDAR) mounted at a fixed position $\mathbf{q}_{s}\in\mathbb{R}^{p}$ and orientation $\mathbf{R}_{s}\in SO(p)$ in the robot’s body frame. This sensor generates noisy distance measurements $\boldsymbol{\eta}(\mathbf{q}_{s},\mathbf{R}_{s})=[\eta_{1}(\mathbf{q}_{s},\mathbf{R}_{s}),...,\eta_{K}(\mathbf{q}_{s},\mathbf{R}_{s})]^{\top}\in[\eta_{\text{min}},\eta_{\text{max}}]^{K}$, where $K$ denotes the number of rays per sensor observation and $\eta_{\text{min}},\eta_{\text{max}}$ denote the sensor’s minimum and maximum range, respectively.
@@
Consider a mobile robot with dynamics as in (1), equipped with a range sensor, operating in an unknown dynamic environment. Design a control policy that safely and efficiently drives the robot to a desired goal position $\mathbf{q}_{G}\in\mathbb{R}^{p}$, while ensuring that the robot’s body satisfies ${\cal B}(\mathbf{x}(t))\subset{\cal F}(t)$ for all $t\geq 0$, despite sensing and environment uncertainty.
@@ 5 System Overview
This section describes the methods we use for localization and mapping, path planning, and control to enable safe robot navigation. Our contribution is the control method, which includes a distributionally robust time-varying control barrier function to guarantee safety in dynamic environments, presented in Sec. 6, and a control Lyapunov function for stable path-following control introduced in Sec. 7. While our control method is applicable to general control-affine systems (1) under certain assumptions, our experiments focus on a wheeled differential-drive robot.
We consider a robot with state $\mathbf{x}:=[x,y,\theta]^{\top}\in{\cal X}\subseteq\mathbb{R}^{2}\times[-\pi,\pi)$, input $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$, and dynamics:
The function $\phi:{\cal X}\rightarrow\mathbb{R}^{2}$ that projects the robot state $\mathbf{x}$ to its position is $\phi(\mathbf{x})=\phi([x,y,\theta]^{\top})=[x,y]^{\top}$. Fig. 2 presents an overview of the robot autonomy components, which are described next.
Localization and Mapping. The robot is equipped with a LiDAR scanner and uses the Hector SLAM algorithm (Kohlbrecher et al., 2011) to estimate its pose and build an occupancy map of the environment. We emphasize that the accuracy and detail of the occupancy map may not be sufficient to ensure safe navigation in a dynamic environment. We use the map for high-level path planning and employ a low-level controller to guarantee safe tracking using CLF and CBF techniques.
@@
Control. Our control approach simultaneously guarantees the robot’s adherence to the planned path $\gamma$ and its safety from collisions in the dynamically changing environment. In Sec. 6, we develop a distributionally robust CBF that uses noisy distance measurements directly to guarantee safety with respect to dynamic obstacles. In Sec. 7, we develop a CLF to track the path $\gamma$ with stability guarantees.
@@ 6 Distributionally Robust Safe Control
We consider a CBF $h(\mathbf{x},t)$ such that its superlevel set ${\cal C}(t)$ in (3) is a subset of $\{\mathbf{x}\in{\cal X}\mid{\cal B}(\mathbf{x}(t))\subset{\cal F}(t)\}$. To develop the DR-CBF formulation, we make the following assumption on the unknown CBF.
@@
The CBF $h(\mathbf{x},t)$ satisfies the following assumptions:
  1. 1.
    $h(\mathbf{x},t)$ has a uniform relative degree of 1 with respect to the system dynamics (1), i.e., the time derivative of $h(\mathbf{x},t)$ along (1) depends explicitly on the control input $\mathbf{u}$;
  2. 2.
    the condition in (5) is satisfied for a linear ${\cal K}_{\infty}$ function $\alpha_{h}$, i.e., $\alpha_{h}(z)=az$ for $a\in\mathbb{R}_{>0}$.
@@
$\text{CBC}(\mathbf{x},\mathbf{u},t)\!=\![\nabla_{\mathbf{x}}h(\mathbf{x},t)]^{\top}\mathbf{F}(\mathbf{x})\underline{\mathbf{u}}+\frac{\partial h(\mathbf{x},t)}{\partial t}+ah(\mathbf{x},t)$
[…]
For each $(\mathbf{x},t)\in{\cal X}\times\mathbb{R}$, we consider the vector $\boldsymbol{\xi}(\mathbf{x},t):=[h(\mathbf{x},t),\nabla_{\mathbf{x}}h(\mathbf{x},t),\frac{\partial h(\mathbf{x},t)}{\partial t}]\in\mathbb{R}^{n+2}$. Since the environment is unknown and the sensor measurements are noisy, $\boldsymbol{\xi}$ cannot be determined exactly. Our objective is to guarantee the satisfaction of constraint (6) despite the uncertainty in $\boldsymbol{\xi}$. To achieve this, we employ distributionally robust optimization techniques, which handle the uncertainty in (6). Before delving into the details of our approach, we review preliminaries of chance constraints and distributionally robust optimization.
@@ 6.1 Chance Constraints and Distributionally Robust Optimization
(Choice of Wasserstein ball radius): There is a connection between the sample size $N$ and the Wasserstein radius $r$ for constructing the ambiguity set ${\cal M}_{N}^{r}$. A distribution $\mathbb{P}$ is light-tailed if there exists an exponent $\rho$ such that $A:=\mathbb{E}_{\mathbb{P}}[\exp{\left\lVert\boldsymbol{\xi}\right\rVert^{\rho}}]=\int_{\Xi}\exp{\left\lVert\boldsymbol{\xi}\right\rVert^{\rho}}\mathbb{P}(d\boldsymbol{\xi})<\infty$. If the true distribution $\mathbb{P}^{*}$ is light-tailed, the choice of $r=r_{N}(\bar{\epsilon})$ given in Esfahani and Kuhn (2018, Theorem 3.5),
@@ 6.2 Distributionally Robust Safety Constraint
At each $(\mathbf{x},t)\in{\cal X}\times\mathbb{R}$, $N$ samples of the vector $\boldsymbol{\xi}=[h(\mathbf{x},t),\nabla_{\mathbf{x}}h(\mathbf{x},t),\frac{\partial h(\mathbf{x},t)}{\partial t}]$ can be obtained, denoted by $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$.
The samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ can be obtained using sensor measurements (we discuss this in detail in Sec. 8.3). Inspired by the CLF-CBF QP formulation in (6), we consider the following distributionally robust formulation to ensure safety with high probability:
$(\mathbf{u}(\mathbf{x},t),\delta)$ $=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{2},\delta\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$
@@
Our previous work (Long et al., 2024) presents a similar result but for a CLF in the context of stabilization under uncertainty. According to Lemma 6.4, the safety of the closed-loop system is guaranteed with high probability. However, the optimization problem in (14) is intractable (Hota et al., 2019; Esfahani and Kuhn, 2018) due to the supremum over the Wasserstein ambiguity set. In Sec. 6.3, we discuss our approach to identify tractable reformulations of (14) and facilitate online safe control synthesis.
@@ 6.3 Tractable Convex Reformulation
Next, demonstrate how the samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ from Assumption 6.3 can be used to obtain a tractable reformulation of (14).
@@
(Distributionally robust safe control synthesis): Given samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ of the barrier $h$, its gradient, and its time derivative, if $(\mathbf{u}^{*},\delta^{*},s^{*},\{\beta_{i}^{*}\}_{i\in[N]})$ is a solution to the quadratic program:
$\min_{\mathbf{u}\in\mathbb{R}^{2},\delta\in\mathbb{R},s\in\mathbb{R},\beta_{i}\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$ (18)
@@
[…]
@@
$rL(\mathbf{u},\mathbf{x})+\inf_{t\in\mathbb{R}}\left[\frac{1}{N}\sum_{i=1}^{N}(-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi}_{i})+s)_{+}-s\epsilon\right]\leq 0$ (21)
where $L(\mathbf{u},\mathbf{x})$ is the Lipschitz constant of $-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi})$ in $\boldsymbol{\xi}$. Now, from (6), we have $\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi})=[a\;\;\mathbf{F}(\mathbf{x})\underline{\mathbf{u}}\;\;1]^{\top}\boldsymbol{\xi}$. Therefore, for each $\mathbf{x}$, we can define the convex function $L:\mathbb{R}^{m}\times{\cal X}\mapsto\mathbb{R}_{>0}$ by
@@
For fixed $(\mathbf{x},\mathbf{u})$, the function $\boldsymbol{\xi}\mapsto-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi})$ is Lipschitz in $\boldsymbol{\xi}$ with constant $L(\mathbf{u},\mathbf{x})$. This is because the Lipschitz constant of a differentiable affine function equals the dual-norm of its gradient, and the dual norm of the $L_{1}$ norm is the $L_{\infty}$ norm. Thus, the following is a conservative approximation of (14),
$\min_{\mathbf{u}\in\mathbb{R}^{2},\delta\in\mathbb{R},s\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$ (23)
@@
$rL($ $\mathbf{u},\mathbf{x})\!+\!\inf_{s\in\mathbb{R}}\left[\frac{1}{N}\sum_{i=1}^{N}(-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi}_{i})+s)_{+}-s\epsilon\right]\!\leq\!0.$
@@
Proposition 6.5 allows control synthesis with distributionally robust safety constraints, only relying on the available samples of $\boldsymbol{\xi}$. The Lipschitz continuity and regularity of distributionally robust controllers are characterized in Mestres et al. (2024). Together with Lemma 6.4, this enables safe robot control with guarantees in unknown dynamic environments without requiring an accurate map reconstruction.

### v2

@@ 4 Problem Formulation
We consider a mobile robot that relies on noisy range measurements to traverse an unknown dynamic environment towards a desired goal. The robot’s motion is governed by control-affine dynamics as in (1). Let $\phi:{\cal X}\rightarrow\mathbb{R}^{p}$ be the mapping that projects the robot’s state $\mathbf{x}$ to its position vector $\phi(\mathbf{x})\in\mathbb{R}^{p}$. We denote the robot orientation as $\mathbf{R}(\mathbf{x})\in\text{SO}(p)$.
Let $\mathcal{B}_{0}\subset\mathbb{R}^{p}$ represent the robot’s shape in its local coordinate frame, i.e., when the robot is at the origin with no rotation or velocity. To describe the shape of the robot in a general state $\mathbf{x}$, we introduce the following linear transformation:
$\mathcal{B}(\mathbf{x})=A(\mathbf{x})\mathcal{B}_{0}+\mathbf{b}(\mathbf{x}),$ (7)
where $A(\mathbf{x})\in\mathbb{R}^{p\times p}$ is an invertible state-dependent transformation matrix (e.g., rotation), and $\mathbf{b}(\mathbf{x})\in\mathbb{R}^{p}$ is a state-dependent translation vector. Both $A$ and $\mathbf{b}$ are assumed to be known continuously differentiable functions of the robot’s state $\mathbf{x}$.
We a signed distance function (SDF) to describe the robot’a shape. Let $\mathcal{S}\subset\mathbb{R}^{p}$ be a closed set. The signed distance function $d:\mathbb{R}^{p}\rightarrow\mathbb{R}$ computes the signed distance from a point $\mathbf{q}\in\mathbb{R}^{p}$ to the set boundary $\partial\mathcal{S}$:
$d_{\mathcal{S}}(\mathbf{q})=\begin{cases}-\min\limits_{\mathbf{q}^{*}\in\partial\mathcal{S}}\|\mathbf{q}-\mathbf{q}^{*}\|,&\text{if }\mathbf{q}\in\mathcal{S},\\ \phantom{-}\min\limits_{\mathbf{q}^{*}\in\partial\mathcal{S}}\|\mathbf{q}-\mathbf{q}^{*}\|,&\text{if }\mathbf{q}\notin\mathcal{S}.\end{cases}$ (8)
This SDF provides a measure of how far a point $\mathbf{q}$ is from the set $\mathcal{S}$. The function takes negative values for points inside $\mathcal{S}$, positive values for points outside, and is zero on the boundary $\partial\mathcal{S}$.
When the set $\mathcal{S}$ represents the robot’s body $\mathcal{B}(\mathbf{x})$, we call its signed distance function the robot SDF. To express the SDF of the robot’s body in terms of its reference shape $\mathcal{B}_{0}$, we apply the linear transformation in (7):
$d(\mathcal{B}(\mathbf{x}),\mathbf{q})=d_{\mathcal{B}_{0}}\left(A(\mathbf{x})^{-1}(\mathbf{q}-\mathbf{b}(\mathbf{x}))\right),$ (9)
where $d_{\mathcal{B}_{0}}(\cdot)$ is the SDF of the reference shape $\mathcal{B}_{0}$. Closed-form robot SDF expressions are available for simple geometries (e.g., circles, spheres) and can be approximated using feed-forward neural networks for more complex geometries (Park et al., 2019; Gropp et al., 2020).
@@
The robot is equipped with a range sensor (e.g., LiDAR) mounted at a fixed position and orientation in the robot’s body frame. The sensor emits multiple rays, each corresponding to a direction in the sensor’s field of view, and generates distance measurements along these rays. The measurements are subject to additive noise, modeled as $\boldsymbol{\eta}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))=\bar{\boldsymbol{\eta}}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))+\mathbf{n}$, where $\bar{\boldsymbol{\eta}}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))$ represents the true distances, and $\mathbf{n}$ is the noise vector, assumed to be bounded and independent for each ray. The resulting noisy measurements are denoted as $\boldsymbol{\eta}(\phi(\mathbf{x}),\mathbf{R}(\mathbf{x}))=[\eta_{1},\ldots,\eta_{K}]^{\top}\in[\eta_{\text{min}},\eta_{\text{max}}]^{K}$, where $K$ denotes the number of rays per sensor observation, and $\eta_{\text{min}},\eta_{\text{max}}$ denote the sensor’s minimum and maximum range, respectively.
The environment includes both static and dynamic obstacles. Let the obstacle set at the initial time $t=0$ be represented as $\mathcal{O}_{0}\subset\mathbb{R}^{p}$. At any time $t$, the obstacle space $\mathcal{O}(t)\subset\mathbb{R}^{p}$ is obtained by applying a smooth transformation to the initial obstacle set:
$\mathcal{O}(t)=A_{\mathcal{O}}(t)\mathcal{O}_{0}+\mathbf{b}_{\mathcal{O}}(t),$ (10)
where $A_{\mathcal{O}}(t)\in\mathbb{R}^{p\times p}$ is a time-dependent transformation matrix and $\mathbf{b}_{\mathcal{O}}(t)\in\mathbb{R}^{p}$ is a time-dependent translation vector, both unknown. To ensure smooth changes in the environment, we assume that both $A_{\mathcal{O}}$ and $\mathbf{b}_{\mathcal{O}}$ are continuously differentiable. This assumption guarantees that obstacles do not experience sudden jumps, expansions, or contractions. Furthermore, obstacles are assumed to have finite thickness, bounded velocities and accelerations, and smooth boundaries to ensure reliable detection.
At time $t$, the boundary of the obstacle space $\partial\mathcal{O}(t)$ is estimated from the noisy sensor measurements $\boldsymbol{\eta}$. The free space, where the robot can operate safely, is the open set $\mathcal{F}(t)=\mathbb{R}^{p}\setminus\mathcal{O}(t)$.
@@
Consider a mobile robot with dynamics given by (1), equipped with a range sensor, operating in an unknown dynamic environment. Design a control policy that drives the robot safely and efficiently to a desired goal position $\mathbf{q}_{G}\in\mathbb{R}^{p}$. The policy must ensure that the robot’s body remains within the free space, i.e., ${\cal B}(\mathbf{x}(t))\subset{\cal F}(t)$ for all $t\geq 0$, while accounting for uncertainties in sensing and state estimation.
@@ 5 System Overview
This section describes the methods we use for localization and mapping, path planning, and control to enable safe robot navigation, Fig. 2 presents an overview of the robot autonomy components. Our contribution is the control method, which includes a distributionally robust time-varying control barrier function to guarantee safety in dynamic environments, presented in Sec. 6, and a control Lyapunov function for stable path-following control introduced in Sec. 7.
Localization and Mapping. The robot is equipped with a LiDAR scanner and uses the Hector SLAM algorithm (Kohlbrecher et al., 2011) to estimate its pose and build an occupancy map of the environment. We emphasize that the efficiency and accuracy of the constructed occupancy map may not be sufficient to ensure safe navigation in a dynamic environment. We use the map for high-level path planning and employ a low-level controller to guarantee safe tracking using CLF and CBF techniques.
@@
Control. Our control approach simultaneously guarantees the robot’s tracking of the planned path $\gamma$ and its safety from collisions in the dynamically changing environment. In Sec. 6, we develop a distributionally robust CBF that uses noisy distance measurements and estimated robot states to directly to guarantee safety with respect to dynamic obstacles. In Sec. 7, we develop a CLF to track the path $\gamma$ with stability guarantees.
@@ 6 Distributionally Robust Safe Control
We consider a CBF $h(\mathbf{x},t)$ with super zero-level set $\mathcal{C}(t)=\{\mathbf{x}\in\mathcal{X}\mid h(\mathbf{x},t)\geq 0\}$ that satisfies
$\mathcal{C}(t)\subseteq\{\mathbf{x}\in\mathcal{X}\mid\mathcal{B}(\mathbf{x}(t))\subseteq\mathcal{F}(t)\},$
where $\mathcal{B}(\mathbf{x}(t))$ represents the robot’s body at state $\mathbf{x}(t)$, and $\mathcal{F}(t)$ is the free space at time $t$. This establishes a connection between the CBF $h$ and the environment geometry. To develop the DR-CBF formulation, we make the following assumption on the unknown CBF.
@@
The CBF $h$ has a uniform relative degree of 1 with respect to the system dynamics (1), i.e., the time derivative of $h(\mathbf{x},t)$ along (1) depends explicitly on the control input $\mathbf{u}$.
@@
$\text{CBC}(\mathbf{x},\mathbf{u},t)=$ (11)
$[\nabla_{\mathbf{x}}h(\mathbf{x},t)]^{\top}\mathbf{F}(\mathbf{x})\underline{\mathbf{u}}+\frac{\partial h(\mathbf{x},t)}{\partial t}+\alpha_{h}(h(\mathbf{x},t))=$
$\underbrace{[\nabla_{\mathbf{x}}h(\mathbf{x},t)^{\top}\mathbf{F}(\mathbf{x}),\;\;\alpha_{h}(h(\mathbf{x},t)),\;\;\frac{\partial h(\mathbf{x},t)}{\partial t}]}_{\boldsymbol{\xi}^{\top}(\mathbf{x},t)}\begin{bmatrix}\underline{\mathbf{u}}\\ 1\\ 1\end{bmatrix}\geq 0,$
where we define an uncertainty vector $\boldsymbol{\xi}(\mathbf{x},t)\in\mathbb{R}^{m+3}$, containing the elements of the time derivative of the barrier function, for each $(\mathbf{x},t)\in{\cal X}\times\mathbb{R}$. Since the environment is unknown and both the sensor measurements and the robot state estimates are noisy, $\boldsymbol{\xi}$ cannot be determined exactly.
Our formulation addresses uncertainties arising from both the barrier function $h$ and the state estimations $\mathbf{x}$. Unlike existing robust or probabilistic CBF approaches, which often focus on uncertainties in system dynamics (Clark, 2019; Long et al., 2022; Dhiman et al., 2023; Wang and Xu, 2023; Breeden and Panagou, 2023; Das and Burdick, 2024), few works tackle the challenges posed by state estimation errors. This stems from the nonlinearity of the dynamics model $\mathbf{F}$ and the barrier function $h$, which makes it difficult to propagate state estimation errors in the control barrier constraint (CBC) in (11).
We address this challenge by leveraging the power of distributionally robust optimization. Instead of explicitly propagating state estimation errors through the system dynamics and barrier functions, we assume access to $M$ state samples $\{\mathbf{x}_{j}\}_{j=1}^{M}$ from a state estimation algorithm. These samples can be obtained, for example, from the Gaussian distributions provided by a Kalman filter (Kalman, 1960), particles generated by a particle filter (Djuric et al., 2003), or a graph-based localization algorithm (Grisetti et al., 2010). The state samples, combined with estimates of $h$ (e.g., obtained directly from the distance measurements $\boldsymbol{\eta}$), are directly used to construct samples of the uncertainty vector $\{\boldsymbol{\xi}_{i}\}_{i=1}^{N}$, as detailed in Sec. 8.3. By managing uncertainties in $\boldsymbol{\xi}$ using distributionally robust optimization, we ensure the satisfaction of constraint (11) without requiring explicit error propagation.
Before introducing the control synthesis formulation, we review the preliminaries of chance constraints and distributionally robust optimization.
@@ 6.1 Chance Constraints and Distributionally Robust Optimization
(Choice of Wasserstein ball radius): There is a connection between the sample size $N$ and the Wasserstein radius $r$ for constructing the ambiguity set ${\cal M}_{N}^{r}$. A distribution $\mathbb{P}$ is light-tailed if there exists an exponent $\rho$ such that $A:=\mathbb{E}_{\mathbb{P}}[\exp{\left\lVert\boldsymbol{\xi}\right\rVert^{\rho}}]=\int_{\Xi}\exp{\left\lVert\boldsymbol{\xi}\right\rVert^{\rho}}\mathbb{P}(d\boldsymbol{\xi})<\infty$. If the true distribution $\mathbb{P}^{*}$ is light-tailed, the choice of $r=r_{N}(\bar{\epsilon})$ given in Esfahani and Kuhn (2018, Theorem 3.5) is
@@ 6.2 Distributionally Robust Safety Constraint
At each $(\mathbf{x},t)\in{\cal X}\times\mathbb{R}$, $N$ samples of the vector $\boldsymbol{\xi}$ in (11) can be obtained, denoted by $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$.
The samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ can be obtained using sensor measurements and state estimation (we discuss this in detail in Sec. 8.3). In many robotic systems, sensing and state estimation may operate at lower frequencies than the control loop. For example, LiDAR measurements or SLAM-based localization may provide updates at about $10$ Hz, while the control loop may require computations at $50$ Hz. Our DR-CBF formulation addresses this challenge by incorporating samples of $\boldsymbol{\xi}$ derived from potentially delayed or uncertain sensor data and state estimations. By accounting for both the asynchrony and uncertainty inherent in sensing and estimation, our approach ensures probabilistic safe performance under realistic conditions.
Inspired by the CLF-CBF QP formulation in (6), we consider the following distributionally robust formulation to ensure safety with high probability:
$(\mathbf{u}(\mathbf{x},t),\delta)$ $=\operatorname*{arg\,min}_{\mathbf{u}\in\mathbb{R}^{m},\delta\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$
@@
Our previous work (Long et al., 2024) presents a similar result but for a CLF in the context of stabilization under uncertainty. According to Lemma 6.4, the safety of the closed-loop system is guaranteed with high probability. However, the optimization problem in (17) is intractable (Hota et al., 2019; Esfahani and Kuhn, 2018) due to the infimum over the Wasserstein ambiguity set. In Sec. 6.3, we discuss our approach to identify tractable reformulations of (17) and facilitate online safe control synthesis.
@@ 6.3 Tractable Convex Reformulation
Next, we demonstrate how the samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ from Assumption 6.3 can be used to obtain a tractable reformulation of (17).
@@
(Distributionally robust safe control synthesis): Given samples $\{\boldsymbol{\xi}_{i}\}_{i\in[N]}$ of $\boldsymbol{\xi}$ as in (11), if $(\mathbf{u}^{*},\delta^{*},s^{*},\{\beta_{i}^{*}\}_{i\in[N]})$ is a solution to the quadratic program:
$\min_{\mathbf{u}\in\mathbb{R}^{m},\delta\in\mathbb{R},s\in\mathbb{R},\beta_{i}\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$ (21)
@@
$r\|\underline{\mathbf{u}}\|_{\infty}\leq s\epsilon-\frac{1}{N}\sum_{i=1}^{N}\beta_{i},$
@@
$rL(\mathbf{u})\!+\!\inf_{s\in\mathbb{R}}\left[\frac{1}{N}\sum_{i=1}^{N}(-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi}_{i})+s)_{+}-s\epsilon\right]\!\leq\!0$ (24)
@@
[…]
The function $\boldsymbol{\xi}\mapsto-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi})$ is Lipschitz in $\boldsymbol{\xi}$ with constant $L(\mathbf{u})$. This is because the Lipschitz constant of a differentiable affine function equals the dual norm of its gradient, and the dual norm of the $L_{1}$ norm is the $L_{\infty}$ norm. Thus, the following is a conservative approximation of (17),
$\min_{\mathbf{u}\in\mathbb{R}^{m},\delta\in\mathbb{R}}\|\mathbf{u}-\mathbf{k}(\mathbf{x})\|^{2}+\lambda\delta^{2},$ (26)
@@
$rL($ $\mathbf{u})\!+\!\inf_{s\in\mathbb{R}}\left[\frac{1}{N}\sum_{i=1}^{N}(-\text{CBC}(\mathbf{x},\mathbf{u},\boldsymbol{\xi}_{i})+s)_{+}-s\epsilon\right]\!\leq\!0.$
@@
Formally, the safe set ${\cal C}(t)$ is defined by a single CBF $h(\mathbf{x},t)$, but its value, gradient, and time derivative (defining $\boldsymbol{\xi}(\mathbf{x},t)$) are not directly known in practice. By leveraging observations of $\boldsymbol{\xi}$ derived from sensor measurements and state estimates, and applying Proposition 6.5, we are able to synthesize safe controllers with distributionally robust guarantees. Due to the convex reformulation of (17b), our approach inherently introduces some conservatism. However, in practice, this conservatism can be effectively managed by tuning the Wasserstein radius $r$ and the safety probability $\epsilon$. These parameters provide flexibility to adapt the formulation based on specific application requirements, sensor capabilities, and state estimation errors.
The Lipschitz continuity and regularity of distributionally robust controllers are characterized in Mestres et al. (2024). Together with Lemma 6.4, this enables safe robot control with point-wise probabilistic guarantees in unknown dynamic environments.
Remark 6.6.
(Uncertainty in System Dynamics): We have assumed that there is no uncertainty in the system dynamics $\mathbf{F}$ to simplify the presentation in Sec. 6.3. However, our approach can be extended if this is not the case as long as samples of $\mathbf{F}(\mathbf{x})$ are available. These samples can be combined with samples of robot state and $h(\mathbf{x},t)$ to construct the uncertainty vector $\{\boldsymbol{\xi}_{i}\}_{i=1}^{N}$ and ensure the validity of Proposition 6.5. $\bullet$